to stop a 90 kg basketball player landing from a jump with a downward vertical How much work is required velocity of 3.48 m/sHow much force is required if the basketball player (Question #2) lands with a hard or soft landing? a For soft landing​​​​​​​
Answer #1

in inche ener chage wandane 0-x90x3.43= -54y.969 J e-549.98 J f =544.968 -fd, =-54498 838.41N f9.0 544.968 d-54y 768 = 2179.8